A random variable X has a Poison distribution with mean equal to 0.4. What is the probability the probability that the random variable is greater than zero.
1
Expert's answer
2020-06-02T20:12:05-0400

For a poisson distribution,

"P(X=x) =\frac {e^\lambda \lambda ^x} {x!}"

"\lambda=0.4"

P(X>0)=1-P(X=0)

"=1-\frac {e^{-0.4} 0.4^0}{0!}"

="1-e^{-0.4}"

=1-0.6703

=0.3297